OpenStudy (anonymous):

if n+3/15=13/45 what is the value of n.

OpenStudy (anonymous):

multiply.

OpenStudy (anonymous):

put the like terms together so you get n= (13/45) - (3/15) = (13/45)-(9/45)= 4/45 :)

OpenStudy (anonymous):

is it